# On the solution of the Collatz problem

### v26

@@ 3.2. Properties of Collatz’s Sequences
Let each Collatz’s Sequence $S_{c}(m_{j},t_{j})$ for $j=0,1,2,\ldots,s-1$ be a sample of the Collatz’s Sequence $S_{c}(m,t)$ that is independent. Then $s$ is the number of independent samples of the Collatz’s Sequence $S_{c}(m,t)$ in the series of Collatz’s Sequences $S_{c}(s)$.
First, by Lemma 3.1, as $s$ tends to infinite, for $j=0,1,2,\ldots,s-1$, there are $s$ even numbers $n_{i_{j}}(m_{j},t_{j})+1=2^{m_{j}+1-i_{j}}3^{i_{j}}t_{j}$.
For another $s/2$ odd numbers $n_{i_{j}}(m_{j},t_{j})=4k+3$, since the even number $n_{i_{j}}(m_{j},t_{j})+1=2^{m_{j}+1-i_{j}}3^{i_{j}}t_{j}=4(k+1)$, there are $12^{m_{j}-1-i_{j}}3^{i_{j}}t_{j}=k+$ 1 and
Then there are $s/2^{2}$ odd numbers $n_{0}(m_{j},t_{j})$ with $m_{j}-i_{j}=1$, $s/2^{3}$ odd numbers $n_{0}(m_{j},t_{j})$ with $m_{j}-i_{j}=2$, and $s/2^{4}$ odd numbers $n_{0}(m_{j},t_{j})$ with $m_{j}-i_{j}=3,\ldots$. Thus as $s$ tends to infinity, i.e., the number of independent samples of the Collatz’s Sequence $S_{c}(m,t)$ tends to infinity, by the law of great numbers, the distribution of the $s$ even numbers $n_{i_{j}}(m_{j},t_{j})+1$ is similar to the distribution of even numbers $n_{e}(\gamma,\lambda)$ stated by Lemma 3.13, so that we have
@@
$=s+\frac{s}{2}\sum_{i=1}^{\infty}i/2^{i}=s+\frac{s}{2}\frac{1/2}{(1-1/2)^{2}}=2s.$
@@
By Lemma 3.5 a $1/2^{r}$ of positive integers $r_{j+1}$ satisfy $r_{j+1}=r$ for $r=1,2,\ldots$. Thus as $s$ tends to infinite, we have
$\sum_{j=0}^{s-1}r_{j+1}=s/2+2s/2^{2}+3s/2^{3}+4s/2^{4}+\cdots$
$=s\sum_{i=1}^{\infty}i/2^{i}=s\frac{1/2}{(1-1/2)^{2}}=2s$
@@
$=\prod_{j=0}^{s-1}(3/5)^{m_{j}+1-i_{j}}\times\prod_{j=0}^{s-1}2^{r_{j+1}}=(3/5)^{2s}2^{2s}>1.$
@@
By the first inequality above for $i=0,1,2,\ldots,m$, the ratio $r_{c}(m,t,i,s)$ of the odd number $n_{i}(m,t)$ to the odd number $n_{i_{s}}(m_{s},t_{s})$ is greater than 1 and does not tend to zero as $s$ tends to infinite, so that there is no Collatz cycle from an odd number $n_{i}(m_{0},t_{0})$ greater than 1 to a same odd number $n_{i_{s}}(m_{s},t_{s})$, except for $n_{i_{s}}(m_{s},t_{s})=n_{0}(m_{0},t_{0})=n_{0}(0,1)=1$, and the odd number $n_{i_{s}}(m_{s},t_{s})$ does not tend to infinite as $s$ tends to infinite. Moreover the inequality $n_{i}(m,t)\geq r_{c}(m,t,0,s)\geq(6/5)^{2s}$ implies that if $s$ tends to infinite then the odd number $n_{i}(m,t)$ must tend to infinite or if the odd number $n_{i}(m,t)$ is finite then the positive integer $s$ must be finite so that the odd number $n_{i_{s}}(m_{s},t_{s})$ is finite and equal to 1 for $n_{i}(m,t)<\infty$.
@@
Hence the odd number $n_{i_{s}}(m_{s},t_{s})$ is finite and equal to 1 as $s$ tends to infinite, and no Collatz cycle contains any odd number greater than 1 in $S_{c}(s)$.
This completes the proof of the lemma. ∎
Lemma 3.13 (Distribution of even numbers).
As a natural number $s_{e}$ tends to infinite, there are $s_{e}/2^{\gamma}$ even numbers $n_{e}(\gamma,\lambda)\leq s_{e}$ for $\gamma=1,2,\ldots$ where
Proof.
Since all even numbers can be expressed as
then as a natural number $s_{e}$ tends to infinite, there are $s_{e}/2^{\gamma}$ even numbers $n_{e}(\gamma,\lambda)\leq s_{e}$ for $\gamma=1,2,\ldots$.
@@ 3.3. Properties of odd numbers $n_{0}(m,t)$ in a range $(N,\rho N]$
Let $n_{0}(m,t)=n\in(N,\rho N]$. Then $n_{0}(m,t)$ in the range $(N,\rho N]$ is equal to $14k+$ 1 so that by Lemma 3.2 there are $m=0$, $t>3$,
@@
Each odd number $n=6x-1$ in the range $(N,\rho N]$ returns to 1.
@@
Since each odd number $n=6x-1$ in the range $(N,\rho N]$ can be rewritten as
@@
Hence each odd number $n=6x-1$ in the range $(N,\rho N]$ returns to 1.
@@
By Lemma 3.5 a half of positive integers $r_{1}$ satisfy $r_{1}=1$ and a quarter of positive integers $r_{1}$ satisfy $r_{1}=2$. Let the notation $P_{r}$ denote the process of mutual references or nested recursions to lemmas 3.17-3.20.
Then the entrance of the process $P_{r}$ is Lemma 3.17, by which $S_{c}(m,t)$ returns to 1 since $S_{c}(m-1,t)$ returns to 1. When the process $P_{r}$ is interrupted or stopped, there must be $r_{1}=1$ in the proof of Lemma 3.17 or $r_{1}=1,2$ in the proof of Lemma 3.20, and the Collatz’s Sequence $S_{c}(m,t)$ returns to 1. Thus the exits of the process $P_{r}$ are Lemma 3.17 and Lemma 3.20 where by Remark 3.21 Lemma 3.20 is equivalent to Lemma 3.17 on the process $P_{r}$.
For each odd number $n_{0}(m,t)=4k+3$ in the range $(N,\rho N]$, let $T_{i}$ denote times of references or recursions to Lemma 3.17 and Lemma 3.20 when the process $P_{r}$ is interrupted or stopped, let $P_{i}(T_{i})$ denote the probabilities of exiting the process $P_{r}$, which are functions of $T_{i}$ and $P_{i}(T_{i})=2^{-T_{i}}$ when $r_{1}=1$ in the proof of Lemma 3.17 and $r_{1}=1,2$ in the proof of Lemma 3.20, and let $P_{t}$ denote the total probability of exiting the process $P_{r}$. Then $P_{t}=\sum_{T_{i}=1}^{T_{h}}P_{i}(T_{i})\to 1$ as $T_{h}\to\infty$, i.e., the process $P_{r}$ is certainly interrupted or stopped. So by Lemma 3.17, the Collatz’s Sequence $S_{c}(m,t)$ returns to 1 and the $n_{0}(m,t)=4k+3$ in the range $(N,\rho N]$ returns to 1.
@@
$=3n_{m}(m-1,t)+2=2(2^{r_{1}-1}3n_{i_{1}}(m_{1},t_{1})+1).$
By Lemma 3.5 a half of positive integers $r_{1}$ satisfy $r_{1}=1$ and a quarter of positive integers $r_{1}$ satisfy $r_{1}=2$.
@@
where $10\leq i_{2}\leq m_{2}-$ 1 and one of the two odd numbers $n_{0}(m_{2}-1,t_{2})$ and $n_{i_{1}}(m_{1},t_{1})$ returns to 1. Then by Lemma 3.18, when one of the two odd numbers $n_{0}(m_{2}-1,t_{2})$ and $n_{i_{1}}(m_{1},t_{1})$ returns to 1 another also returns to 1, and by Lemma 3.17, when one of the two Collatz’s Sequences $S_{c}(m_{2},t_{2})$ and $S_{c}(m_{2}-1,t_{2})$ returns to 1 another also returns to 1, so that when one of the two Collatz’s Sequences $S_{c}(m,t)$ and $S_{c}(m-1,t)$ returns to 1 another also returns to 1.
@@
By Lemma 3.22, when one of the two numbers
$13\times 4^{r}n_{i_{1}}(m_{1},t_{1})+1\texttt{ and }3^{r+1}n_{i_{1}}(m_{1},t_{1})+$ 1
returns to 1 another also returns to 1 where $0\leq i_{1}\leq m_{1}$. So when one of the two odd numbers $n_{m+1}(m,t)$ and $13^{r+1}n_{i_{1}}(m_{1},t_{1})+$ 1 returns to 1 another also returns to 1. Let $n=3^{r}n_{i_{1}}(m_{1},t_{1})$ where one of the two odd numbers $n$ and $n_{i_{1}}(m_{1},t_{1})$ returns to 1. Then by Lemma 3.18, when one of the two odd numbers $n$ and $n_{i_{1}}(m_{1},t_{1})$ returns to 1 another also returns to 1, so that when one of the two Collatz’s Sequences $S_{c}(m,t)$ and $S_{c}(m-1,t)$ returns to 1 another also returns to 1.
For example, when $n_{i_{1}}(m_{1},t_{1})$ returns to 1, i.e., $S_{c}(m-1,t)$ returns to 1, then by Lemma 3.18 the odd number $n$ returns to 1, i.e., the number $13n+1=3^{r+1}n_{i_{1}}(m_{1},t_{1})+$ 1 returns to 1, and by Lemma 3.22 the odd number $13\times 4^{r}n_{i_{1}}(m_{1},t_{1})+$ 1 returns to 1, so that $S_{c}(m,t)$ returns to 1.
@@
By Lemma 3.23, when one of the two odd numbers
$13\times 4^{r}2n_{i_{1}}(m_{1},t_{1})+1\texttt{ and }3^{r+1}2n_{i_{1}}(m_{1},t_{1})+$ 1
returns to 1 another also returns to 1. So when one of the two odd numbers $n_{m+1}(m,t)$ and $13^{r+1}2n_{i_{1}}(m_{1},t_{1})+$ 1 returns to 1 another also returns to 1. Let
@@
where $0\leq i<m_{4}$ and one of the two odd numbers $n_{i}(m_{4}-1,t_{4})$ and $n_{i_{1}}(m_{1},t_{1})$ returns to 1. Then by Lemma 3.18, when one of the two odd numbers $n_{i}(m_{4}-1,t_{4})$ and $n_{i_{1}}(m_{1},t_{1})$ returns to 1 another also returns to 1, and by Lemma 3.17, when one of the two Collatz’s Sequences $S_{c}(m_{4},t_{4})$ and $S_{c}(m_{4}-1,t_{4})$ returns to 1 another also returns to 1, so that when one of the two Collatz’s Sequences $S_{c}(m_{4},t_{4})$ and $S_{c}(m-1,t)$ returns to 1 another also returns to 1 or when one of the two Collatz’s Sequences $S_{c}(m,t)$ and $S_{c}(m-1,t)$ returns to 1 another also returns to 1.
For example, when $n_{i_{1}}(m_{1},t_{1})$ returns to 1, i.e., $S_{c}(m-1,t)$ returns to 1, then by Lemma 3.18 the odd number $n_{i}(m_{4}-1,t_{4})$ returns to 1, and by Lemma 3.17 $S_{c}(m_{4},t_{4})$ returns to 1, then by Lemma 3.23 the odd number $13\times 4^{r}2n_{i_{1}}(m_{1},t_{1})+$ 1 returns to 1, so that the Collatz’s Sequence $S_{c}(m,t)$ returns to 1.
@@
Given an odd number $n=3^{r}n_{i}(m,t)$ where $r>0$ and $i\leq m$, when one of the two odd numbers $n$ and $n_{i}(m,t)$ returns to 1 another also returns to 1.
@@
Let an odd number $n_{i_{1}}(m_{1},t_{1})=n_{i}(m,t)$. Then when one of the two odd numbers $n_{i_{1}}(m_{1},t_{1})$ and $n_{i}(m,t)$ returns to 1 another also returns to 1.
Suppose that when one of the two odd numbers $n_{i_{1}}(m_{1},t_{1})=3^{r-1}n_{i}(m,t)$ and $n_{i}(m,t)$ returns to 1 another also returns to 1 where $0\leq i_{1}\leq m_{1}$ and $r>0$. Then there are
@@
By Lemma 3.17 when one of the two odd numbers $n_{i_{1}+1}(m_{1},t_{1})$ and $n_{i_{1}+1}(m_{1}-1,t_{1})$ returns to 1 another also returns to 1, so that when one of the two odd numbers $n$ and $n_{i_{1}+1}(m_{1}-1,t_{1})$ returns to 1 another also returns to 1, and when one of the two odd numbers $n$ and $n_{i}(m,t)$ returns to 1 another also returns to 1.
@@
Otherwise for $i_{1}+1=m_{1}$, $n_{i_{1}+1}(m_{1}-1,t_{1})=2^{r_{2}}n_{i_{2}}(m_{2},t_{2})$ where $r_{2}\geq 1$, $0\leq i_{2}\leq m_{2}$, and $n_{i_{2}}(m_{2},t_{2})$ is odd and not dividable by $3$. By Lemma 3.17 when one of the two odd numbers $n_{i_{1}+1}(m_{1},t_{1})$ and $n_{i_{1}+1}(m_{1}-1,t_{1})$ returns to 1 another also returns to 1, so that when one of the two odd numbers $n_{i_{1}+1}(m_{1},t_{1})$ and $n_{i_{2}}(m_{2},t_{2})$ returns to 1 another also returns to 1. Then there are
@@
By Lemma 3.19 when one of the two odd numbers $13\times 2^{r_{2}}n_{i_{2}}(m_{2},t_{2})+$ 1 and $n_{i_{2}}(m_{2},t_{2})$ returns to 1 another also returns to 1, so that when one of the two odd numbers $n$ and $n_{i}(m,t)$ returns to 1 another also returns to 1.
For example, when $n_{i}(m,t)$ returns to 1 $S_{c}(m_{1},t_{1})$ returns to 1, then by Lemma 3.17 $S_{c}(m_{1}-1,t_{1})$ and $S_{c}(m_{2},t_{2})$ return to 1, and by Lemma 3.19 the odd number $13\times 2^{r_{2}}n_{i_{2}}(m_{2},t_{2})+$ 1 return to 1, so that $n$ returns to 1.
@@
$13^{r_{2}+1}n_{i_{2}}(m_{2},t_{2})=2^{r_{4}}n_{i_{4}}(m_{4},t_{4})+$ 1 .
By Lemma 3.18 when one of the two odd numbers $3^{r_{2}+1}n_{i_{2}}(m_{2},t_{2})$ and $n_{i_{2}}(m_{2},t_{2})$ returns to 1 another also returns to 1.
By Lemma 3.20 when one of the two odd numbers $12^{r_{4}}n_{i_{4}}(m_{4},t_{4})+$ 1 and $n_{i_{4}}(m_{4},t_{4})$ returns to 1 another also returns to 1, so that when one of the two odd numbers $n_{i_{2}}(m_{2},t_{2})$ and $n_{i_{4}}(m_{4},t_{4})$ returns to 1 another also returns to 1 or when one of the two odd numbers $n_{i_{2}}(m_{2},t_{2})$ and $n_{0}(m_{3},t_{3})$ returns to 1 another also returns to 1, i.e., when one of the two odd numbers $n$ and $n_{i}(m,t)$ returns to 1 another also returns to 1.
For example, when $n_{i}(m,t)$ returns to 1 $S_{c}(m_{1},t_{1})$ and $S_{c}(m_{2},t_{2})$ return to 1, then by Lemma 3.18 the odd number $3^{r_{2}+1}n_{i_{2}}(m_{2},t_{2})$ that equals to $12^{r_{4}}n_{i_{4}}(m_{4},t_{4})+$ 1 returns to 1, and by Lemma 3.20 $S_{c}(m_{4},t_{4})$ and $S_{c}(m_{3},t_{3})$ return to 1, so that $n$ returns to 1.
@@
Given an odd number $n=3\times 2^{r_{1}}n_{i_{1}}(m_{1},t_{1})+1$ where $r_{1}>0$, $0\leq i_{1}\leq m_{1}$, and $n_{i_{1}}(m_{1},t_{1})$ is odd and not dividable by 3, when one of the two odd numbers $n$ and $n_{i_{1}}(m_{1},t_{1})$ returns to 1 another also returns to 1.
@@
where $10\leq i\leq m-$ 1 and one of the two odd numbers $n_{i}(m-1,t)$ and $n_{i_{1}}(m_{1},t_{1})$ returns to 1. Then by Lemma 3.18, when one of the two Collatz’s Sequences $S_{c}(m-1,t)$ and $S_{c}(m_{1},t_{1})$ returns to 1 another also returns to 1, and by Lemma 3.17, when one of the two Collatz’s Sequences $S_{c}(m,t)$ and $S_{c}(m-1,t)$ returns to 1 another also returns to 1, so that when one of the two odd numbers $n$ and $n_{i_{1}}(m_{1},t_{1})$ returns to 1 another also returns to 1.

### v32

@@ 3.2. Properties of Collatz’s Sequences
First, by Lemma 3.13, $1/2^{k}$ of positive integers $m_{j}-i_{j}+1$ satisfy $m_{j}-i_{j}+1=k$ for $k=1,2,\ldots$ as $s$ tends to infinity, so that we have
@@
$s+\frac{s}{2}\sum_{i=1}^{\infty}i/2^{i}=s+\frac{s}{2}\frac{1/2}{(1-1/2)^{2}}=2s.$
@@
By Lemma 3.5 $1/2^{r}$ of positive integers $r_{j+1}$ satisfy $r_{j+1}=r$ for $r=1,2,\ldots$. Thus as $s$ tends to infinite, we have
$\sum_{j=0}^{s-1}r_{j+1}=s/2+2s/2^{2}+3s/2^{3}+4s/2^{4}+\cdots=s\sum_{i=1}^{\infty}i/2^{i}=s\frac{1/2}{(1-1/2)^{2}}=2s$
@@
$\prod_{j=0}^{s-1}(3/5)^{m_{j}+1-i_{j}}\times\prod_{j=0}^{s-1}2^{r_{j+1}}=(3/5)^{2s}2^{2s}>1.$
@@
By the first inequality above for $i=0,1,2,\ldots,m$, the ratio $r_{c}(m,t,i,s)$ of the odd number $n_{i}(m,t)$ to the odd number $n_{i_{s}}(m_{s},t_{s})$ is greater than 1 and does not tend to zero as $s$ tends to infinite, so that there is no nontrivial Collatz cycle from an odd number $n_{i}(m_{0},t_{0})$ greater than 1 to a same odd number $n_{i_{s}}(m_{s},t_{s})$, except for $n_{i_{s}}(m_{s},t_{s})=n_{0}(m_{0},t_{0})=n_{0}(0,1)=1$, and the odd number $n_{i_{s}}(m_{s},t_{s})$ does not tend to infinite as $s$ tends to infinite. Moreover the inequality $n_{i}(m,t)\geq r_{c}(m,t,0,s)\geq(6/5)^{2s}$ implies that if $s$ tends to infinite then the odd number $n_{i}(m,t)$ must tend to infinite or if the odd number $n_{i}(m,t)$ is finite then the positive integer $s$ must be finite so that the odd number $n_{i_{s}}(m_{s},t_{s})$ is equal to 1 for $n_{i}(m,t)<\infty$.
@@
Hence no nontrivial Collatz cycle exists in $S_{c}(s)$ and the odd number $n_{i_{s}}(m_{s},t_{s})$ in $S_{c}(s)$ is equal to 1 as $s$ tends to infinite.
@@ 3.3. Properties of odd numbers $n_{0}(m,t)$ in a range $(N,\rho N]$
Let $n_{0}(m,t)=n\in(N,\rho N]$. Then $n_{0}(m,t)$ in the range $(N,\rho N]$ is equal to $4k+1$ so that by Lemma 3.2 there are $m=0$, $t>3$,
@@
Each odd number n=6x1n=6x-1 in the range $(N,\rho N]$ returns to 1.
@@
Since each odd number n=6x1n=6x-1 in the range $(N,\rho N]$ can be rewritten as
@@
Hence each odd number n=6x1n=6x-1 in the range $(N,\rho N]$ returns to 1.
@@
By Lemma 3.5 one half of positive integers $r_{1}$ satisfy $r_{1}=1$ and one quarter of positive integers $r_{1}$ satisfy $r_{1}=2$. Let the notation $P_{r}$ denote the process of mutual references or nested recursions to lemmas 3.22-3.25.
Then the entrance of the process $P_{r}$ is Lemma 3.22, by which $S_{c}(m,t)$ returns to 1 since $S_{c}(m-1,t)$ returns to 1. When the process $P_{r}$ is interrupted or stopped, the Collatz’s Sequence $S_{c}(m,t)$ returns to 1 and there must be $r_{1}=1$ in the proof of Lemma 3.22 or $r_{1}=1,2$ in the proof of Lemma 3.25. Thus the exits of the process $P_{r}$ are Lemma 3.22 and Lemma 3.25 where by Remark 3.26 Lemma 3.25 is equivalent to Lemma 3.22 on the process $P_{r}$.
For each odd number $n_{0}(m,t)=4k+3$ in the range $(N,\rho N]$, let $T_{i}$ denote times of references or recursions to Lemma 3.22 and Lemma 3.25 when the process $P_{r}$ is interrupted or stopped, let $P_{i}(T_{i})$ denote the probabilities of exiting the process $P_{r}$ when $r_{1}=1$ in the proof of Lemma 3.22 and $r_{1}=1,2$ in the proof of Lemma 3.25, which are functions of $T_{i}$ and $P_{i}(T_{i})=2^{-T_{i}}$, and let $P_{t}$ denote the total probability of exiting the process $P_{r}$. Then $P_{t}=\sum_{T_{i}=1}^{T_{h}}P_{i}(T_{i})\to 1$ as $T_{h}\to\infty$, i.e., the process $P_{r}$ is certainly interrupted or stopped. So by Lemma 3.22, the Collatz’s Sequence $S_{c}(m,t)$ returns to 1 and the $n_{0}(m,t)=4k+3$ in the range $(N,\rho N]$ returns to 1.
@@
$3n_{m}(m-1,t)+2=2(2^{r_{1}-1}3n_{i_{1}}(m_{1},t_{1})+1).$
By Lemma 3.5 one half of positive integers $r_{1}$ satisfy $r_{1}=1$ and one quarter of positive integers $r_{1}$ satisfy $r_{1}=2$.
@@
where $0\leq i_{2}\leq m_{2}-1$ and one of two numbers $n_{0}(m_{2}-1,t_{2})$ and $n_{i_{1}}(m_{1},t_{1})$ returns to 1. Then by Lemma 3.23, when one of two numbers $n_{0}(m_{2}-1,t_{2})$ and $n_{i_{1}}(m_{1},t_{1})$ returns to 1 another also returns to 1, and by Lemma 3.22, when one of the two Collatz’s Sequences $S_{c}(m_{2},t_{2})$ and $S_{c}(m_{2}-1,t_{2})$ returns to 1 another also returns to 1, so that when one of the two Collatz’s Sequences $S_{c}(m,t)$ and $S_{c}(m-1,t)$ returns to 1 another also returns to 1.
@@
By Lemma 3.27, when one of two numbers
$3\times 4^{r}n_{i_{1}}(m_{1},t_{1})+1\texttt{ and }3^{r+1}n_{i_{1}}(m_{1},t_{1})+1$
returns to 1 another also returns to 1 where $0\leq i_{1}\leq m_{1}$. So when one of two numbers $n_{m+1}(m,t)$ and $3^{r+1}n_{i_{1}}(m_{1},t_{1})+1$ returns to 1 another also returns to 1. Let $n=3^{r}n_{i_{1}}(m_{1},t_{1})$ where one of two numbers $n$ and $n_{i_{1}}(m_{1},t_{1})$ returns to 1. Then by Lemma 3.23, when one of two numbers $n$ and $n_{i_{1}}(m_{1},t_{1})$ returns to 1 another also returns to 1, so that when one of the two Collatz’s Sequences $S_{c}(m,t)$ and $S_{c}(m-1,t)$ returns to 1 another also returns to 1.
For example, when $n_{i_{1}}(m_{1},t_{1})$ returns to 1, i.e., $S_{c}(m-1,t)$ returns to 1, then by Lemma 3.23 the odd number $n$ returns to 1, i.e., the number $3n+1=3^{r+1}n_{i_{1}}(m_{1},t_{1})+1$ returns to 1, and by Lemma 3.27 the odd number $3\times 4^{r}n_{i_{1}}(m_{1},t_{1})+1$ returns to 1, so that $S_{c}(m,t)$ returns to 1.
@@
By Lemma 3.28, when one of two numbers
$3\times 4^{r}2n_{i_{1}}(m_{1},t_{1})+1\texttt{ and }3^{r+1}2n_{i_{1}}(m_{1},t_{1})+1$
returns to 1 another also returns to 1. So when one of two numbers $n_{m+1}(m,t)$ and $3^{r+1}2n_{i_{1}}(m_{1},t_{1})+1$ returns to 1 another also returns to 1. Let
@@
where $0\leq i<m_{4}$ and one of two numbers $n_{i}(m_{4}-1,t_{4})$ and $n_{i_{1}}(m_{1},t_{1})$ returns to 1. Then by Lemma 3.23, when one of two numbers $n_{i}(m_{4}-1,t_{4})$ and $n_{i_{1}}(m_{1},t_{1})$ returns to 1 another also returns to 1, and by Lemma 3.22, when one of the two Collatz’s Sequences $S_{c}(m_{4},t_{4})$ and $S_{c}(m_{4}-1,t_{4})$ returns to 1 another also returns to 1, so that when one of the two Collatz’s Sequences $S_{c}(m_{4},t_{4})$ and $S_{c}(m-1,t)$ returns to 1 another also returns to 1 or when one of the two Collatz’s Sequences $S_{c}(m,t)$ and $S_{c}(m-1,t)$ returns to 1 another also returns to 1.
For example, when $n_{i_{1}}(m_{1},t_{1})$ returns to 1, i.e., $S_{c}(m-1,t)$ returns to 1, then by Lemma 3.23 the odd number $n_{i}(m_{4}-1,t_{4})$ returns to 1, and by Lemma 3.22 $S_{c}(m_{4},t_{4})$ returns to 1, then by Lemma 3.28 the odd number $3\times 4^{r}2n_{i_{1}}(m_{1},t_{1})+1$ returns to 1, so that the Collatz’s Sequence $S_{c}(m,t)$ returns to 1.
@@
Given an odd number $n=3^{r}n_{i}(m,t)$ where $r>0$ and $i\leq m$, when one of two numbers $n$ and $n_{i}(m,t)$ returns to 1 another also returns to 1.
@@
Let an odd number $n_{i_{1}}(m_{1},t_{1})=n_{i}(m,t)$. Then when one of two numbers $n_{i_{1}}(m_{1},t_{1})$ and $n_{i}(m,t)$ returns to 1 another also returns to 1.
Suppose that when one of two numbers $n_{i_{1}}(m_{1},t_{1})=3^{r-1}n_{i}(m,t)$ and $n_{i}(m,t)$ returns to 1 another also returns to 1 where $0\leq i_{1}\leq m_{1}$ and $r>0$. Then there are
@@
By Lemma 3.22 when one of two numbers $n_{i_{1}+1}(m_{1},t_{1})$ and $n_{i_{1}+1}(m_{1}-1,t_{1})$ returns to 1 another also returns to 1, so that when one of two numbers $n$ and $n_{i_{1}+1}(m_{1}-1,t_{1})$ returns to 1 another also returns to 1, and when one of two numbers $n$ and $n_{i}(m,t)$ returns to 1 another also returns to 1.
@@
Otherwise for $i_{1}+1=m_{1}$, $n_{i_{1}+1}(m_{1}-1,t_{1})=2^{r_{2}}n_{i_{2}}(m_{2},t_{2})$ where $r_{2}\geq 1$, $0\leq i_{2}\leq m_{2}$, and $n_{i_{2}}(m_{2},t_{2})$ is odd and not dividable by $3$. By Lemma 3.22 when one of two numbers $n_{i_{1}+1}(m_{1},t_{1})$ and $n_{i_{1}+1}(m_{1}-1,t_{1})$ returns to 1 another also returns to 1, so that when one of two numbers $n_{i_{1}+1}(m_{1},t_{1})$ and $n_{i_{2}}(m_{2},t_{2})$ returns to 1 another also returns to 1. Then there are
@@
By Lemma 3.24 when one of two numbers $3\times 2^{r_{2}}n_{i_{2}}(m_{2},t_{2})+1$ and $n_{i_{2}}(m_{2},t_{2})$ returns to 1 another also returns to 1, so that when one of two numbers $n$ and $n_{i}(m,t)$ returns to 1 another also returns to 1.
For example, when $n_{i}(m,t)$ returns to 1 $S_{c}(m_{1},t_{1})$ returns to 1, then by Lemma 3.22 $S_{c}(m_{1}-1,t_{1})$ and $S_{c}(m_{2},t_{2})$ return to 1, and by Lemma 3.24 the odd number $3\times 2^{r_{2}}n_{i_{2}}(m_{2},t_{2})+1$ return to 1, so that $n$ returns to 1.
@@
$3^{r_{2}+1}n_{i_{2}}(m_{2},t_{2})=2^{r_{4}}n_{i_{4}}(m_{4},t_{4})+1.$
By Lemma 3.23 when one of two numbers $3^{r_{2}+1}n_{i_{2}}(m_{2},t_{2})$ and $n_{i_{2}}(m_{2},t_{2})$ returns to 1 another also returns to 1.
By Lemma 3.25 when one of two numbers $2^{r_{4}}n_{i_{4}}(m_{4},t_{4})+1$ and $n_{i_{4}}(m_{4},t_{4})$ returns to 1 another also returns to 1, so that when one of two numbers $n_{i_{2}}(m_{2},t_{2})$ and $n_{i_{4}}(m_{4},t_{4})$ returns to 1 another also returns to 1 or when one of two numbers $n_{i_{2}}(m_{2},t_{2})$ and $n_{0}(m_{3},t_{3})$ returns to 1 another also returns to 1, i.e., when one of two numbers $n$ and $n_{i}(m,t)$ returns to 1 another also returns to 1.
For example, when $n_{i}(m,t)$ returns to 1 $S_{c}(m_{1},t_{1})$ and $S_{c}(m_{2},t_{2})$ return to 1, then by Lemma 3.23 the odd number $3^{r_{2}+1}n_{i_{2}}(m_{2},t_{2})$ that equals to $2^{r_{4}}n_{i_{4}}(m_{4},t_{4})+1$ returns to 1, and by Lemma 3.25 $S_{c}(m_{4},t_{4})$ and $S_{c}(m_{3},t_{3})$ return to 1, so that $n$ returns to 1.
@@
Given an odd number $n=3\times 2^{r_{1}}n_{i_{1}}(m_{1},t_{1})+1$ where $r_{1}>0$, $0\leq i_{1}\leq m_{1}$, and $n_{i_{1}}(m_{1},t_{1})$ is odd and not dividable by 3, when one of two numbers $n$ and $n_{i_{1}}(m_{1},t_{1})$ returns to 1 another also returns to 1.
@@
where $0\leq i\leq m-1$ and one of two numbers $n_{i}(m-1,t)$ and $n_{i_{1}}(m_{1},t_{1})$ returns to 1. Then by Lemma 3.23, when one of the two Collatz’s Sequences $S_{c}(m-1,t)$ and $S_{c}(m_{1},t_{1})$ returns to 1 another also returns to 1, and by Lemma 3.22, when one of the two Collatz’s Sequences $S_{c}(m,t)$ and $S_{c}(m-1,t)$ returns to 1 another also returns to 1, so that when one of two numbers $n$ and $n_{i_{1}}(m_{1},t_{1})$ returns to 1 another also returns to 1.